OpenStudy (megannicole51):

evaluate:

OpenStudy (megannicole51):

\[\int\limits_{0}^{1} \int\limits_{e ^{y}}^{e} \frac{ 1+e ^{x} }{ \ln(x) }dxdy\]

OpenStudy (megannicole51):

@agent0smith

hartnn (hartnn):

so are you taught change of order ?

OpenStudy (agent0smith):

\[\Large \int\limits\limits_{0}^{1} \int\limits\limits_{e ^{y}}^{e} \frac{ 1+e ^{x} }{ \ln(x) }dxdy\] cos it's easier to read

OpenStudy (megannicole51):

\[\int\limits_{e ^{y}}^{e}\frac{ 1+e ^{y} }{ \ln(x) }dx\]

OpenStudy (megannicole51):

first?

OpenStudy (megannicole51):

no bc my professor is incompetent

OpenStudy (megannicole51):

please teach me

hartnn (hartnn):

yeah, thats the problem, that function cannot be integrated thats where change of order comes into picture

OpenStudy (megannicole51):

how do u know that

OpenStudy (agent0smith):

\[\Large \int\limits\limits_{e ^{y}}^{e}\frac{ 1+e ^{y} }{ \ln(x) }dx\]above it was a 1+e^x now it's 1+e^y...?

hartnn (hartnn):

wait what ?

OpenStudy (megannicole51):

its x sorry i messed up

hartnn (hartnn):

ok, we can use change of order when 1) the function to be integrated is very difficult to integrate 2) or its not possible to integrate in standard functions we'll know this only by trying it out

hartnn (hartnn):

if you want to make sure, try it out by whichever formulas you know. if you find it difficult/impossible , you'll then appreciate the use of change of order

OpenStudy (primeralph):

Why bother? Just use the bounds to bind the integral.

hartnn (hartnn):

the first step for changing the order will be to find the region the limits are x = e^y to x =e can you plot these 2 curves ? (x= e will be a line)

OpenStudy (megannicole51):

theres not an easier way to do this? on my exam i wont be able to plot the curves bc i cant do it offhand

hartnn (hartnn):

if change of order is involved, getting the region of integration is must.

OpenStudy (megannicole51):

lovely:(

OpenStudy (primeralph):

Or get a really cool calculator.

OpenStudy (megannicole51):

no calculators can be used for college exams.

OpenStudy (megannicole51):

i never use mine anymore

OpenStudy (isaiah.feynman):

Your brain will bud. :)

OpenStudy (megannicole51):

okay im not trying to be mean but my calc final is monday. @hartnn can we get back to the problem i really just need to know how to do it step by step.

hartnn (hartnn):

you can plot x= e^y by hand ?

hartnn (hartnn):

x= e is easy,

OpenStudy (megannicole51):

no i can't but i can look it up

OpenStudy (dan815):